OpenStudy (nannie):

Please show me how to work this. 11z-66/4÷z-6/6z

OpenStudy (anonymous):

z(11z-66)/4 -6/6z = (11z^2 - 66)/4 - 1/z) = 2.75z^2 - 16.5 - 1/z=